In a study of distances traveled by buses before the first major engine failure, a sample of 191 buses resulted in a mean of 96,700 miles and a standard deviation of 37,500 miles. At the 0.05 level of signıficance, test the manufacturer's claim that the mean distance traveled before a major engine failure is more than 90,000 miles.


1. Claim:

Ho:

Ha:


2. Level of Significance:

Test- statistic:

Tails in Distribution:


3. Reject Ho if:


4. Compute for the value of the test statistics.


5. Make a decision:


6. State the conclusion in terms of the original problem


1
Expert's answer
2021-03-30T11:09:37-0400

"H_0" : the mean distance traveled before a major engine failure is more than 90,000 miles

"H_a" : the mean distance traveled before a major engine failure is not more than 90,000 miles

Level of Significance: "\alpha=0.05"

Test- statistic:

"z=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\frac{96700-90000}{37500/\sqrt{191}}=2.45"

p-value:

"p=P(z>2.45)=1-0.9929=0.0071"

"p<\alpha\implies" we have to reject "H_0"

Conclusion: the mean distance traveled before a major engine failure is not more than 90,000 miles.